OpenStudy (anonymous):

integrate 1/(x^2+6x+18)

OpenStudy (anonymous):

partial fraction decomp... see attached

OpenStudy (anonymous):

oops, may be something different. just a sec.

OpenStudy (anonymous):

i think it is partial fractions but I have no idea what to do with the denomenator

OpenStudy (anonymous):

it's an irreducible quadratic

OpenStudy (anonymous):

i think you have to complete the square... hold on

OpenStudy (anonymous):

try this

OpenStudy (anonymous):

\[x^2+6x+18 = (x+3)^2 + 9\] let u = x+3 then you can use an inverse trig fucntion to integrate \[\int\limits \frac{ 1 }{ x^2+6x+18 }\,dx=\int\limits \frac{ 1 }{ \left( x+3 \right)^2+9}\,dx=\int\limits \frac{ 1 }{ u^2+9}\,du\]

OpenStudy (anonymous):

then can i just use the formula \[\int\limits dx/(x^2+a^2) = 1/a \arctan (x/a) + C\]

OpenStudy (anonymous):

wrong trig function... have a look

OpenStudy (anonymous):

my bad... you got it.

OpenStudy (anonymous):

be careful how you write in here, though. better to use \frac then just \ 1/n arctan(n/a) or \(\large{\frac{1}{n} \arctan \left(\frac{n}{a}\right)}\)

OpenStudy (anonymous):

i thought you had the arctan on the bottom

OpenStudy (anonymous):

you good to go?

OpenStudy (anonymous):

yes, thanks a lot

OpenStudy (anonymous):

you're welcome!